OpenStudy (jackellyn):

y' +y =2 y(0)=2 Find a solution using integrating factor.

OpenStudy (jackellyn):

The integrating factor would be e^x right?

OpenStudy (anonymous):

2(y - 4x^2)dx + (x)dy = 0 2(y - 4x^2) + (x)(dy/dx) = 0 2y - 8x^2 + xy' = 0 xy' + 2y = 8x^2 y' + (2/x)y = 8x I.F. = e^ int(2/x)dx = e^(2lnx) = e^(lnx^2) = x^2 (x^2)y' + (x^2)(2/x)y = (x^2)8x (x^2)y' + (2x)y = 8x^3 d((x^2)y)/dx = 8x^3 d((x^2)y) = 8x^3 dx (x^2)y = 2x^4 + C

OpenStudy (cwrw238):

IF = e^(INT1dx) = e^x right

OpenStudy (jackellyn):

What would my next step be @cwrw238 ?

OpenStudy (jackellyn):

thanks @abayomi12 , it might be better if you explained but you did help me with another question i had so thumbs up for you (not literally)

OpenStudy (cwrw238):

next step s to multiply each term by e^x

OpenStudy (jackellyn):

\[e ^{x}y' +e ^{x}y = 2e^{x}\]

OpenStudy (cwrw238):

ok - note that left hand side is the result of differentiating ye^x so: ye^x = INT 2e^x

OpenStudy (jackellyn):

\[y=e ^{-x}(2e ^{x}+C)\]

OpenStudy (jackellyn):

\[y=2+Ce ^{-x}\] is it -x or x?

OpenStudy (jackellyn):

so y=2, awesome thanks @cwrw238

OpenStudy (cwrw238):

-x

OpenStudy (cwrw238):

yw